### v1

@@ 19 Initialize $V_{free}(x,t),V_{free}(x,t)$
20 $V_{free}(x,t)=\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}(\phi_{free}(x^{\prime},t^{\prime})\times V(x^{\prime},t^{\prime})/\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\phi_{free}(x^{\prime},t^{\prime})$
21 $V_{cong}(x,t)=\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}(\phi_{cong}(x^{\prime},t^{\prime})\times V(x^{\prime},t^{\prime})/\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\phi_{cong}(x^{\prime},t^{\prime})$
22 $V_{min}(x,t)=min(V_{free}(x,t),V_{cong}(x,t))$
23 $tanh(x,t)=tanh(\nicefrac{{(V_{c}-V_{min}(x,t))}}{{\Delta V}})$
24 $w=0.5\times(1+tahnh(x,t)$

### v3

@@ 19 Initialize $V_{free}(x,t),V_{free}(x,t)$
20 $V_{free}(x,t)=\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\Bigl{(}\phi_{free}(x^{\prime},t^{\prime})\times V(x^{\prime},t^{\prime})\Bigr{)}\Big{/}\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\phi_{free}(x^{\prime},t^{\prime})$
21 $V_{cong}(x,t)=\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\Bigl{(}\phi_{cong}(x^{\prime},t^{\prime})\times V(x^{\prime},t^{\prime})\Bigr{)}\Big{/}\sum_{x^{\prime}=x-\nicefrac{{\sigma}}{{2}}}^{x+\nicefrac{{\sigma}}{{2}}}\phi_{cong}(x^{\prime},t^{\prime})$
22 $V_{min}(x,t)=\min\Bigl{(}V_{free}(x,t),V_{cong}(x,t)\Bigr{)}$
23 $tanh(x,t)=\tanh\Bigl{(}\nicefrac{{(V_{c}-V_{min}(x,t))}}{{\Delta V}}\Bigr{)}$
24 $w=0.5\times\Bigl{(}1+\tanh(x,t)\Bigr{)}$